OpenStudy (anonymous):

find the 5th,10th,and 100th terms of each sequence . -19,-22,-25,-28

OpenStudy (anonymous):

this is an arithmetic sequence

OpenStudy (anonymous):

a1 = -19 what about d ? can you find d by yourself ?

OpenStudy (anonymous):

yes it is i know the formula but i need help on how do i get -19 to -22

OpenStudy (anonymous):

wait i got it -3

OpenStudy (anonymous):

d = -22 - (-19) (since d = a(n+1) - an)

OpenStudy (anonymous):

very good!

OpenStudy (anonymous):

so now you can find the formula for the nth term an

OpenStudy (anonymous):

and then it's done..

OpenStudy (anonymous):

ok i got it now thanks:)

OpenStudy (anonymous):

yw